# Intersection and Union Hierarchies of Deterministic

### v2

@@ 6.1 Grammars and Grammatical Trees
A (formal) grammar is a quadruple $G=\langle N,\Sigma,P,S\rangle$, where $N$ is a finite set of terminal symbols (or terminals), $\Sigma$ is a finite set of non-terminal symbols (or nonterminals), $P$ is a set of production rules (or productions), and $S$ is the start symbol. We use the notation $\Rightarrow$ to denote a derivation and $\Rightarrow_{r}$ to indicate a rightmost derivation, in which any production should be first applied to the rightmost nonterminal. The notation $\Rightarrow_{r}^{*}$ indicates the reflexive, transitive closure of $\Rightarrow_{r}$. Given any nonterminal $A\in N$ and any string $\in(N\cup\Sigma)^{*}, ww$ is called a right canonical grammatical form of $A$ if $A\Rightarrow_{r}^{*}w$. A grammar $G$ is said to be reduced if (i) there is no reduction of the form $A\to A$ for any $A\in N$, (ii) for any $A\in N$, there is a string $w\in\Sigma^{*}$ satisfying $A\Rightarrow^{+}w$, and (iii) for any $A\in N$, there are strings $x,y\in(N\cup\Sigma)^{*}$ for which $S\Rightarrow^{*}xAy$.
@@
Associated with a grammar $G$, we deal with a labeled rooted tree $T$ with a set $D$ of nodes and the root $r$ as well as two binary relations $\triangleright$ and $\prec$ and a labeling function $\rho:D\to N\cup\Sigma\cup\{\varepsilon\}$. We further expand $\rho$ to a function from $D^{*}$ to $(V\cup\Sigma)^{*}$ by setting $\rho(v_{1},v_{2},\ldots,v_{m})=\rho(v_{1})\rho(v_{2})\cdots\rho(v_{m})$ for all nodes $v_{1},v_{2},\ldots,v_{m}\in D$, where $D^{*}$ denotes the set of all finite sequences of nodes in $D$. We say that a production of the form $A\to\sigma_{1}\sigma_{2}\cdots\sigma_{m}$ (resp., $A\to\varepsilon$) for $A\in N$ and $\sigma_{1},\sigma_{2},\ldots,\sigma_{m}\in N\cup\Sigma$ expresses an elementary subtree $T$ if the root of $T$ has label $A$ and all children have the labels $\sigma_{1},\sigma_{2},\ldots,\sigma_{m}$ from left to right (resp., have label $\varepsilon$). Such a tree $T$ is called a grammatical (derivation) tree of $G$ if (i) for every elementary subtree $T^{\prime}$ of $T$, there is a production in $P$ that expresses $T^{\prime}$ and (ii) the string $\rho(leaf(T))$ belongs to $\Sigma^{*}$. Such a grammatical tree is expressed as $(D,\triangleright,\prec,r,\rho)$. When the root $r$ satisfies $\rho(r)=S$, $T$ is called a derivation tree of $G$. If $T$ further satisfies $\rho(r)\Rightarrow^{*}_{r}\rho(leaf(T))$ in $G$, then $T$ is called right canonical. It is often useful to relax the above condition (ii) to the following condition: (ii) $\rho(leaf(T))\in(N\cup\Sigma)^{*}$. With this new condition (ii) together with (i), we call $T$ a partial grammatical tree of $G$. Let ${\cal T}_{G}$ denote the set of all partial grammatical trees of $G$. We say that $T$ is extendable to grammatical trees if there exists a rightmost canonical grammatical tree $T^{\prime}$ of $G$ such that $T$ is a subtree of $T^{\prime}$ with the same root.
@@
Let $v_{1}$ and $v_{2}$ denote two nodes of $T$ with $v_{1}\leadsto v_{2}$ and let $\alpha$ and $\beta$ be two sequences of leaves of $T$. We say that $(v_{1},v_{2})$ generates $(\alpha,\beta)$ in $T$ if $leaf(T(v_{1}))=\alpha\cdot leaf(T(v_{2}))\cdot\beta$. Here, the dot “ $\cdot$ ” indicates a concatenation of two sequences. For example, $(v_{1},v_{2},v_{3})\cdot(w_{1},w_{2})$ indicates $(v_{1},v_{2},v_{3},w_{1},w_{2})$. Notice that $\alpha$ or $\beta$ (or both) in $(\alpha,\beta)$ may be empty. Consider a path $p$ from the root $r$ to a leaf in $T$. Given two distinct nodes $v_{1}$ and $v_{2}$ lying on the path $p$ with $v_{1}\leadsto v_{2}$, we define $st[v_{1},v_{2}]$ to be the string pair $(\rho(\alpha),\rho(\beta))$ obtained from a pair $(\alpha,\beta)$ that is generated by $(v_{1},v_{2})$ in $T$. Moreover, let $ST[p]$ be the set $\{st[v_{1},v_{2}]\mid v_{1},v_{2}\in p,v_{1}\leadsto v_{2},v_{1}\neq v_{2}\}$.
@@
Figure 4 illustrates two right canonical partial grammatical trees.
@@
Take an index $j\in[1,i_{1}]_{\mathbb{Z}}$ satisfying that $\rho_{1}(c_{j})=\varepsilon$ and the production $\rho_{1}(u_{j}^{(1)})\rightarrow\rho_{1}(c_{j})$ (i.e., $B^{\prime}_{j}\rightarrow\varepsilon$) is the most recently applied one while generating $T_{1}$. Since $T_{1}$ is right canonical, the production $\rho_{1}(u_{j}^{(1)})\rightarrow\varepsilon$ cannot be applied earlier than the production $\rho_{1}(v_{1}^{(1)})\rightarrow\rho_{1}(c_{i_{1}-2},c_{i_{1}-1},c_{i_{1}})$ (i.e., $B_{1}\rightarrow\beta^{\prime}$). Even after the application of $B_{1}\rightarrow\beta^{\prime}$, we cannot apply $B^{\prime}_{j}\rightarrow\varepsilon$ because $\rho_{1}(c_{i_{1}})\in N$. Hence, there is no chance of applying $B^{\prime}_{j}\rightarrow\varepsilon$. This is a contradiction. Therefore, $\rho_{1}(c_{j})$ cannot be $\varepsilon$. ∎
@@
Let us look into $T_{2}$. Since $d_{i_{2}}$ is the rightmost leaf labeled with a nonterminal, we conclude that $\rho_{2}(d_{j})=\varepsilon$ for all indices $j\in[i_{2}+1,t]_{\mathbb{Z}}$. This means that, for each index $j\in[i_{2}+1,t]_{\mathbb{Z}}$, there is a node $u^{(2)}_{j}$ such that $d_{j}$ is the only child of $u^{(2)}_{j}$. For brevity, we write $B^{\prime\prime}_{j}$ for $\rho_{2}(u^{(2)}_{j})$. Its associated production $\rho_{2}(u^{(2)}_{j})\rightarrow\rho_{2}(d_{j})$ is $B^{\prime\prime}_{j}\rightarrow\varepsilon$. We delete $d_{i_{2}+1}$ from $T_{2}$ and call the obtained subtree by $T^{\prime}_{2}$. From this $T^{\prime}_{2}$, we obtain $S\Rightarrow_{r}^{*}\rho_{2}(d_{1},\ldots,d_{i_{2}})\rho_{2}(u^{(2)}_{i_{2}+1})\rho_{2}(d_{i_{2}+2},\ldots,d_{t})\rho_{2}(d_{t+1},\ldots,d_{n})=\alpha^{\prime\prime}\beta^{\prime\prime}B^{\prime}_{i_{2}+1}az\Rightarrow_{r}\alpha^{\prime\prime}\beta^{\prime\prime}az$. By the LR(1) property of $G$, it then follows that $B_{1}=B^{\prime\prime}_{i_{2}+1}$ and $\beta^{\prime}=\varepsilon$. This is a contradiction because of $\beta^{\prime}=\rho_{1}(c_{s-2},c_{s-1},c_{s})$. Therefore, $i_{2}=t$ must follow. ∎
Figure 5: Cases: (i) and (ii)
@@
(i) Consider the case where $(c_{i_{1}-2},c_{i_{1}-1},c_{i_{1}})$ and $(d_{i_{2}-2},d_{i_{2}-1},d_{i_{2}})$ are children of $v_{1}^{(1)}$ and $v_{1}^{(2)}$, respectively. Since $i_{1}=s$, $\rho_{1}(c_{i_{1}+1})=a$ holds. We delete all children of $v_{1}^{(1)}$ from $T_{1}$ and all children of $v_{1}^{(2)}$ from $T_{2}$. We respectively call the obtained subtrees by $T^{\prime\prime}_{1}$ and $T^{\prime\prime}_{2}$. Note that $leaf(T^{\prime\prime}_{1})=(c_{1},\ldots,c_{i_{1}-3},v_{1}^{(1)},c_{i_{1}+1},\ldots,c_{m})$ with $|leaf(T^{\prime\prime}_{1})|+2=|leaf(T_{1})|$. We then obtain the derivations $S\Rightarrow_{r}^{*}\alpha^{\prime}\rho_{1}(v_{1}^{(1)})ay=\alpha^{\prime}B_{1}ay\Rightarrow_{r}\alpha^{\prime}\beta^{\prime}ay$ by applying the production $B_{1}\rightarrow\beta^{\prime}$. Similarly, since $leaf(T^{\prime\prime}_{2})=(d_{1},\ldots,d_{i_{2}-3},v_{1}^{(2)},d_{i_{2}+1},\ldots,d_{n})$ with $|leaf(T^{\prime\prime}_{2})|+2=|leaf(T_{2})|$, we obtain $S\Rightarrow_{r}^{*}\alpha^{\prime\prime}\rho_{2}(v_{1}^{(2)})az=\alpha^{\prime\prime}B_{2}az\Rightarrow_{r}\alpha^{\prime\prime}\beta^{\prime\prime}az$ by applying $B_{2}\rightarrow\beta^{\prime\prime}$. The LR(1) property of $G$ then yields both $B_{1}=B_{2}$ and $\beta^{\prime}=\beta^{\prime\prime}$.
@@
Let $F_{1}$ and $F_{2}$ respectively denote the single node trees consisting of $v_{1}^{(1)}$ and $v_{1}^{(2)}$ and let $F^{\prime}_{1}$ and $F^{\prime}_{2}$ respectively denote the subtrees of $T_{1}$ and $T_{2}$, which are rooted at $v_{1}^{(1)}$ and $v_{1}^{(2)}$. Let $^{(s-2)}T^{\prime\prime}_{1}=(F_{11},\ldots,F_{1l_{1}})$ and $^{(t-2)}T^{\prime\prime}_{2}=(F_{21},\ldots,F_{2l_{2}})$. By the definitions of $T^{\prime\prime}_{1}$ and $T^{\prime\prime}_{2}$, it follows that $F_{1l_{1}}=F_{1}$ and $F_{2l_{2}}=F_{2}$. Therefore, $T^{\prime\prime}_{1}$ and $T^{\prime\prime}_{2}$ satisfy the conditions (1)–(5) of the lemma. Since $|leaf(T^{\prime\prime}_{1})|<|leaf(T_{1})|$ and $|leaf(T^{\prime\prime}_{2})|<|leaf(T_{2})|$, we apply induction hypothesis to $T^{\prime\prime}_{1}$ and $T^{\prime\prime}_{2}$ and then obtain s2=t2s-2=t-2 and $^{(s-2)}T^{\prime\prime}_{1}\equiv{}^{(t-2)}T^{\prime\prime}_{2}$. It thus follows that $l_{1}=l_{2}$ and $F_{1i}\equiv F_{2i}$ for all $i\in[l_{1}]$. By replacing $F_{1}$ and $F_{2}$ respectively with $F^{\prime}_{1}$ and $F^{\prime}_{2}$, we obtain $^{(s)}T_{1}=(F_{11},\ldots,F_{1l_{1}-1},F^{\prime}_{1})$ and $^{(t)}T_{2}=(F_{21},\ldots,F_{2l_{2}-1},F^{\prime}_{2})$. Since $F^{\prime}_{1}\equiv F^{\prime}_{2}$ and $l_{1}=l_{2}$, we conclude that $^{(s)}T_{1}\equiv{}^{(t)}T_{2}$.
@@
(ii) Consider the case where the children of $v_{1}^{(1)}$ are $(c_{i_{1}-1},c_{i_{1}},v_{0}^{(1)})$ with $\rho_{1}(c_{i_{1}+1})=a$ and the children of $v_{1}^{(2)}$ are $(d_{i_{2}-1},d_{i_{2}},v_{0}^{(2)})$. Let $X=\rho_{1}(c_{i_{1}})$. There are two grammatical trees $\tilde{T}_{1}$ and $\tilde{T}_{2}$ for which their roots have label $X$ and all leaves of them are labeled by terminal symbols. We then expand $T_{1}$ and $T_{2}$ by attaching $\tilde{T}_{1}$ and $\tilde{T}_{2}$ to their leaves $c_{i_{1}}$ and $d_{i_{2}}$, respectively, and we call the resulting trees by $T_{1}^{(ext)}$ and $T_{2}^{(ext)}$. We set $s^{\prime}=i_{1}-1$ and $t^{\prime}=i_{2}-1$. The newly constructed trees $T_{1}^{(ext)}$ and $T_{2}^{(ext)}$ satisfy the conditions (1)–(6) of the lemma.
@@
(iii) We study the case where $(c_{i_{1}-2},c_{i_{1}-1},c_{i_{1}})$ are the children of $v_{1}^{(1)}$ but $(d_{i_{2}-1},d_{i_{2}},v_{0}^{(2)})$ are the children of $v_{1}^{(2)}$. In this case, we obtain the derivations $S\Rightarrow_{r}^{*}\rho_{1}(c_{1},\ldots,c_{i_{1}-3})\rho_{1}(v_{1}^{(1)})\rho_{1}(c_{i_{1}+1},\ldots,c_{m})=\alpha^{\prime}B_{1}ay\Rightarrow_{r}\alpha^{\prime}\beta^{\prime}ay$ by applying $B_{1}\rightarrow\beta^{\prime}$ and also $S\Rightarrow_{r}^{*}\rho_{2}(d_{1},\ldots,d_{i_{2}})\rho_{2}(v_{0}^{(2)})\rho_{2}(d_{i_{2}+2},\ldots,d_{n})=\alpha^{\prime\prime}\beta^{\prime\prime}A_{2}z\Rightarrow_{r}\alpha^{\prime\prime}\beta^{\prime}az$ by applying the production $A_{2}\rightarrow a$. The LR(1) property of $G$ then concludes that $B_{1}=A_{2}$ and $\beta^{\prime}=a$. This is a contradiction with $|\beta^{\prime}|=3$. Therefore, this case never happens.
@@
(2) We next examine the second case of $i_{1}<s$. We set $p=s-i_{1}$ so that $c_{i_{1}+p}=c_{s}$. This implies that $\rho_{1}(c_{i_{1}+1},\ldots,c_{s})=\varepsilon$. Let $u^{(1)}_{j}$ denote the parent of $c_{j}$. Notice that $c_{j}$ is the only child of $u^{(2)}_{j}$. Similarly, we set $p^{\prime}=t-i_{2}$. By Claim 3 (by swapping $T_{1}$ and $T_{2}$), $p^{\prime}$ cannot be 0, and thus we obtain $p^{\prime}>0$. We also write $u^{(2)}_{j}$ for the parent of $d_{j}$. We delete $c_{i_{1}}$ from $T_{1}$ and $d_{i_{2}+1}$ from $T_{2}$ call the resulting subtrees by $T^{\prime}_{1}$ and $T^{\prime}_{2}$, respectively. For simplicity, let $A^{\prime}_{j}=\rho_{1}(u^{(1)}_{j})$ and $A^{\prime\prime}_{j}=\rho_{2}(u^{(2)}_{j})$.
@@
Consider an $\mathrm{LR}(1)$ grammar $G=\langle N,\Sigma,P,S\rangle$ for $L$ and assume that $G$ already becomes reduced. As noted in Section 6.1, we further assume without loss of generality that $G$ is in Greibach normal form. Let $k=|N|$ and let $h$ denote the maximum length of the right-hand side of any production of $G$. We define $c=2^{k\log{h}}(k+1)(h-1)$.
@@
Let us first consider the basis case of $d=1$. Take two arbitrary strings $w=xy$ and $w^{\prime}=xz$ in $L$ with $|x|,|y|,|z|\geq 1$ and $|x|>c$. Let $a$ denote the last symbol of $x$ and let $x^{(-)}$ denote the string obtained from $x$ by deleting its last symbol. Note that $x=x^{(-)}a$ and $|x^{(-)}|\geq c$.
@@
We split $xy$ to two parts: a prefix $\hat{x}$ and a suffix $\hat{y}$ of $xy$ satisfying $\hat{x}\hat{y}=xy$. Let $\hat{a}$ denote the first symbol of $\hat{y}$ if $\hat{y}\neq\varepsilon$. Take a path $p_{i}$ from the root $r_{1}$ to $c_{s_{i}+1}$ with $\rho_{1}(c_{s_{i}+1})=\hat{a}$.

### v3

@@ 6.1 Grammars and Grammatical Trees
A (formal) grammar is a quadruple $G=\langle N,\Sigma,P,S\rangle$, where $N$ is a finite set of terminal symbols (or terminals, for short), $\Sigma$ is a finite set of non-terminal symbols (or nonterminals), $P$ is a set of production rules (or productions), and $S$ is the start symbol. We use the notation $\Rightarrow$ to denote a derivation and $\Rightarrow_{r}$ to indicate a rightmost derivation, in which any production should be first applied to the rightmost nonterminal. The notation $\Rightarrow_{r}^{*}$ indicates the reflexive, transitive closure of $\Rightarrow_{r}$. Given any nonterminal $A\in N$ and any string $\in(N\cup\Sigma)^{*}, ww$ is called a right canonical grammatical form of $A$ if $A\Rightarrow_{r}^{*}w$. A grammar $G$ is said to be reduced if (i) there is no reduction of the form $A\to A$ for any $A\in N$, (ii) for any $A\in N$, there is a string $w\in\Sigma^{*}$ satisfying $A\Rightarrow^{+}w$, and (iii) for any $A\in N$, there are strings $x,y\in(N\cup\Sigma)^{*}$ for which $S\Rightarrow^{*}xAy$.
@@
Associated with a grammar $G$, we deal with a labeled rooted tree $T$ with a set $D$ of nodes, the root $r$ as well as two binary relations $\triangleright$ and $\prec$, and a labeling function $\rho:D\to N\cup\Sigma\cup\{\varepsilon\}$. We further expand $\rho$ to a function from $D^{*}$ to $(V\cup\Sigma)^{*}$ by setting $\rho(v_{1},v_{2},\ldots,v_{m})=\rho(v_{1})\rho(v_{2})\cdots\rho(v_{m})$ for all nodes $v_{1},v_{2},\ldots,v_{m}\in D$, where $D^{*}$ denotes the set of all finite sequences of nodes in $D$. We say that a production of the form $A\to\sigma_{1}\sigma_{2}\cdots\sigma_{m}$ (resp., $A\to\varepsilon$) for $A\in N$ and $\sigma_{1},\sigma_{2},\ldots,\sigma_{m}\in N\cup\Sigma$ expresses an elementary subtree $T$ for which the root of $T$ has label $A$ and all children have the labels $\sigma_{1},\sigma_{2},\ldots,\sigma_{m}$ from left to right (resp., have label $\varepsilon$). Such a tree $T$ is called a grammatical (derivation) tree of $G$ if (i) for every elementary subtree $T^{\prime}$ of $T$, there is a production in $P$ that expresses $T^{\prime}$ and (ii) the string $\rho(leaf(T))$ belongs to $\Sigma^{*}$. Such a grammatical tree is expressed as $(D,\triangleright,\prec,r,\rho)$. When the root $r$ satisfies $\rho(r)=S$, $T$ is called a derivation tree of $G$. If $T$ further satisfies $\rho(r)\Rightarrow^{*}_{r}\rho(leaf(T))$ in $G$, then $T$ is called right canonical. It is often useful to relax the above condition (ii) to the following condition: (ii) $\rho(leaf(T))\in(N\cup\Sigma)^{*}$. With this new condition (ii) together with (i), we call $T$ a partial grammatical tree of $G$. Let ${\cal T}_{G}$ denote the set of all partial grammatical trees of $G$. We say that $T$ is extendable to grammatical trees if there exists a rightmost canonical grammatical tree $T^{\prime}$ of $G$ such that $T$ is a subtree of $T^{\prime}$ with the same root.
@@
Let $v_{1}$ and $v_{2}$ denote two nodes of $T$ with $v_{1}\leadsto v_{2}$ and let $\alpha$ and $\beta$ be two sequences of leaves of $T$. We say that $(v_{1},v_{2})$ generates $(\alpha,\beta)$ in $T$ if $leaf(T(v_{1}))=\alpha\cdot leaf(T(v_{2}))\cdot\beta$. Here, the dot “ $\cdot$ ” indicates the concatenation of two sequences. For example, $(v_{1},v_{2},v_{3})\cdot(w_{1},w_{2})$ indicates $(v_{1},v_{2},v_{3},w_{1},w_{2})$. Notice that $\alpha$ or $\beta$ (or both) in $(\alpha,\beta)$ may be empty. Consider a path $p$ from the root $r$ to a leaf in $T$. Given two distinct nodes $v_{1}$ and $v_{2}$ lying on the path $p$ with $v_{1}\leadsto v_{2}$, we define $st[v_{1},v_{2}]$ to be the string pair $(\rho(\alpha),\rho(\beta))$ obtained from a pair $(\alpha,\beta)$ that is generated by $(v_{1},v_{2})$ in $T$. Moreover, let $ST[p]$ be the set $\{st[v_{1},v_{2}]\mid v_{1},v_{2}\in p,v_{1}\leadsto v_{2},v_{1}\neq v_{2}\}$.
@@
Figure 4 illustrates two right canonical partial grammatical trees discussed in Lemma 6.1.
@@
Take an index $j\in[1,i_{1}]_{\mathbb{Z}}$ satisfying that $\rho_{1}(c_{j})=\varepsilon$ and that the production $\rho_{1}(u_{j}^{(1)})\rightarrow\rho_{1}(c_{j})$ (i.e., $B^{\prime}_{j}\rightarrow\varepsilon$) is the most recently applied one while generating $T_{1}$. Since $T_{1}$ is right canonical, the production $\rho_{1}(u_{j}^{(1)})\rightarrow\varepsilon$ cannot be applied earlier than the production $\rho_{1}(v_{1}^{(1)})\rightarrow\rho_{1}(c_{i_{1}-2},c_{i_{1}-1},c_{i_{1}})$ (i.e., $B_{1}\rightarrow\beta^{\prime}$). Even after the application of $B_{1}\rightarrow\beta^{\prime}$, we cannot apply $B^{\prime}_{j}\rightarrow\varepsilon$ because $\rho_{1}(c_{i_{1}})\in N$. Hence, there is no chance of applying $B^{\prime}_{j}\rightarrow\varepsilon$. This is a contradiction. Therefore, $\rho_{1}(c_{j})$ cannot be $\varepsilon$. ∎
@@
Let us look into $T_{2}$. Since $d_{i_{2}}$ is the rightmost leaf labeled with a nonterminal, we conclude that $\rho_{2}(d_{j})=\varepsilon$ for all indices $j\in[i_{2}+1,t]_{\mathbb{Z}}$. This means that, for each index $j\in[i_{2}+1,t]_{\mathbb{Z}}$, there is a node $u^{(2)}_{j}$ for which $d_{j}$ is the only child of $u^{(2)}_{j}$. For brevity, we write $B^{\prime\prime}_{j}$ for $\rho_{2}(u^{(2)}_{j})$. Its associated production $\rho_{2}(u^{(2)}_{j})\rightarrow\rho_{2}(d_{j})$ is $B^{\prime\prime}_{j}\rightarrow\varepsilon$. We delete $d_{i_{2}+1}$ from $T_{2}$ and call the obtained subtree by $T^{\prime}_{2}$. From this $T^{\prime}_{2}$, we obtain $S\Rightarrow_{r}^{*}\rho_{2}(d_{1},\ldots,d_{i_{2}})\rho_{2}(u^{(2)}_{i_{2}+1})\rho_{2}(d_{i_{2}+2},\ldots,d_{t})\rho_{2}(d_{t+1},\ldots,d_{n})=\alpha^{\prime\prime}\beta^{\prime\prime}B^{\prime}_{i_{2}+1}az\Rightarrow_{r}\alpha^{\prime\prime}\beta^{\prime\prime}az$. By the LR(1) property of $G$, it then follows that $B_{1}=B^{\prime\prime}_{i_{2}+1}$ and $\beta^{\prime}=\varepsilon$. This is a contradiction because of $\beta^{\prime}=\rho_{1}(c_{s-2},c_{s-1},c_{s})$. Therefore, $i_{2}=t$ must follow. ∎
Figure 5: Graphical descriptions of the cases (i) and (ii)
@@
(i) Consider the case where $(c_{i_{1}-2},c_{i_{1}-1},c_{i_{1}})$ and $(d_{i_{2}-2},d_{i_{2}-1},d_{i_{2}})$ are children of $v_{1}^{(1)}$ and $v_{1}^{(2)}$, respectively. See Fig. 5(i). Since $i_{1}=s$, $\rho_{1}(c_{i_{1}+1})=a$ holds. We delete all children of $v_{1}^{(1)}$ from $T_{1}$ and all children of $v_{1}^{(2)}$ from $T_{2}$. We respectively call the obtained subtrees by $T^{\prime\prime}_{1}$ and $T^{\prime\prime}_{2}$. Note that $leaf(T^{\prime\prime}_{1})=(c_{1},\ldots,c_{i_{1}-3},v_{1}^{(1)},c_{i_{1}+1},\ldots,c_{m})$ with $|leaf(T^{\prime\prime}_{1})|+2=|leaf(T_{1})|$. We then obtain the derivations $S\Rightarrow_{r}^{*}\alpha^{\prime}\rho_{1}(v_{1}^{(1)})ay=\alpha^{\prime}B_{1}ay\Rightarrow_{r}\alpha^{\prime}\beta^{\prime}ay$ by applying the production $B_{1}\rightarrow\beta^{\prime}$. Similarly, since $leaf(T^{\prime\prime}_{2})=(d_{1},\ldots,d_{i_{2}-3},v_{1}^{(2)},d_{i_{2}+1},\ldots,d_{n})$ with $|leaf(T^{\prime\prime}_{2})|+2=|leaf(T_{2})|$, we obtain $S\Rightarrow_{r}^{*}\alpha^{\prime\prime}\rho_{2}(v_{1}^{(2)})az=\alpha^{\prime\prime}B_{2}az\Rightarrow_{r}\alpha^{\prime\prime}\beta^{\prime\prime}az$ by applying $B_{2}\rightarrow\beta^{\prime\prime}$. The LR(1) property of $G$ then yields both $B_{1}=B_{2}$ and $\beta^{\prime}=\beta^{\prime\prime}$.
@@
Let $F_{1}$ and $F_{2}$ respectively denote the single-node trees consisting of $v_{1}^{(1)}$ and $v_{1}^{(2)}$ and let $F^{\prime}_{1}$ and $F^{\prime}_{2}$ respectively denote the subtrees of $T_{1}$ and $T_{2}$, which are rooted at $v_{1}^{(1)}$ and $v_{1}^{(2)}$. Let $^{(s-2)}T^{\prime\prime}_{1}=(F_{11},\ldots,F_{1l_{1}})$ and $^{(t-2)}T^{\prime\prime}_{2}=(F_{21},\ldots,F_{2l_{2}})$. By the definitions of $T^{\prime\prime}_{1}$ and $T^{\prime\prime}_{2}$, it follows that $F_{1l_{1}}=F_{1}$ and $F_{2l_{2}}=F_{2}$. Therefore, $T^{\prime\prime}_{1}$ and $T^{\prime\prime}_{2}$ satisfy the conditions (1)–(5) of the lemma. Since $|leaf(T^{\prime\prime}_{1})|<|leaf(T_{1})|$ and $|leaf(T^{\prime\prime}_{2})|<|leaf(T_{2})|$, we apply induction hypothesis to $T^{\prime\prime}_{1}$ and $T^{\prime\prime}_{2}$ and then obtain s2=t2s-2=t-2 and $^{(s-2)}T^{\prime\prime}_{1}\equiv{}^{(t-2)}T^{\prime\prime}_{2}$. It thus follows that $l_{1}=l_{2}$ and $F_{1i}\equiv F_{2i}$ for all $i\in[l_{1}]$. By replacing $F_{1}$ and $F_{2}$ respectively with $F^{\prime}_{1}$ and $F^{\prime}_{2}$, we obtain $^{(s)}T_{1}=(F_{11},\ldots,F_{1l_{1}-1},F^{\prime}_{1})$ and $^{(t)}T_{2}=(F_{21},\ldots,F_{2l_{2}-1},F^{\prime}_{2})$. Since $F^{\prime}_{1}\equiv F^{\prime}_{2}$ and $l_{1}=l_{2}$, we conclude that $^{(s)}T_{1}\equiv{}^{(t)}T_{2}$.
@@
(ii) Consider the case where the children of $v_{1}^{(1)}$ are $(c_{i_{1}-1},c_{i_{1}},v_{0}^{(1)})$ with $\rho_{1}(c_{i_{1}+1})=a$ and the children of $v_{1}^{(2)}$ are $(d_{i_{2}-1},d_{i_{2}},v_{0}^{(2)})$. See Fig. 5(ii). Let $X=\rho_{1}(c_{i_{1}})$. There are two grammatical trees $\tilde{T}_{1}$ and $\tilde{T}_{2}$ for which their roots have label $X$ and all leaves of them are labeled by terminal symbols. We then expand $T_{1}$ and $T_{2}$ by attaching $\tilde{T}_{1}$ and $\tilde{T}_{2}$ to their leaves $c_{i_{1}}$ and $d_{i_{2}}$, respectively, and we call the resulting trees by $T_{1}^{(ext)}$ and $T_{2}^{(ext)}$. We set $s^{\prime}=i_{1}-1$ and $t^{\prime}=i_{2}-1$. The newly constructed trees $T_{1}^{(ext)}$ and $T_{2}^{(ext)}$ satisfy the conditions (1)–(6) of the lemma.
@@
(iii) We study the case where $(c_{i_{1}-2},c_{i_{1}-1},c_{i_{1}})$ are the children of $v_{1}^{(1)}$ but $(d_{i_{2}-1},d_{i_{2}},v_{0}^{(2)})$ are the children of $v_{1}^{(2)}$. In this case, we obtain the derivations $S\Rightarrow_{r}^{*}\rho_{1}(c_{1},\ldots,c_{i_{1}-3})\rho_{1}(v_{1}^{(1)})\rho_{1}(c_{i_{1}+1},\ldots,c_{m})=\alpha^{\prime}B_{1}ay\Rightarrow_{r}\alpha^{\prime}\beta^{\prime}ay$ by applying the production $B_{1}\rightarrow\beta^{\prime}$ and also $S\Rightarrow_{r}^{*}\rho_{2}(d_{1},\ldots,d_{i_{2}})\rho_{2}(v_{0}^{(2)})\rho_{2}(d_{i_{2}+2},\ldots,d_{n})=\alpha^{\prime\prime}\beta^{\prime\prime}A_{2}z\Rightarrow_{r}\alpha^{\prime\prime}\beta^{\prime}az$ by applying $A_{2}\rightarrow a$. The LR(1) property of $G$ then concludes that $B_{1}=A_{2}$ and $\beta^{\prime}=a$. This is a contradiction with $|\beta^{\prime}|=3$. Therefore, this case never happens.
@@
(2) We next examine the second case of $i_{1}<s$. We set $p=s-i_{1}$ so that $c_{i_{1}+p}=c_{s}$. This implies that $\rho_{1}(c_{i_{1}+1},\ldots,c_{s})=\varepsilon$. Let $u^{(1)}_{j}$ denote the parent of $c_{j}$. Notice that $c_{j}$ is the only child of $u^{(1)}_{j}$. Similarly, we set $p^{\prime}=t-i_{2}$. By Claim 3 (by swapping $T_{1}$ and $T_{2}$), $p^{\prime}$ cannot be 0, and thus we obtain $p^{\prime}>0$. We also write $u^{(2)}_{j}$ for the parent of $d_{j}$. We delete $c_{i_{1}}$ from $T_{1}$ and $d_{i_{2}+1}$ from $T_{2}$ and we then call the resulting subtrees by $T^{\prime}_{1}$ and $T^{\prime}_{2}$, respectively. For simplicity, let $A^{\prime}_{j}=\rho_{1}(u^{(1)}_{j})$ and $A^{\prime\prime}_{j}=\rho_{2}(u^{(2)}_{j})$.
@@
Let us first consider the basis case of $d=1$. Consider an $\mathrm{LR}(1)$ grammar $G=\langle N,\Sigma,P,S\rangle$ for $L$ and assume that $G$ already becomes reduced. As noted in Section 6.1, we further assume without loss of generality that $G$ is in Greibach normal form. Let $k=|N|$ and let $h$ denote the maximum length of the right-hand side of any production of $G$. We define $c=2^{k\log{h}}(k+1)(h-1)$.
Take two arbitrary strings $w=xy$ and $w^{\prime}=xz$ in $L$ with $|x|,|y|,|z|\geq 1$ and $|x|>c$. Let $a$ denote the last symbol of $x$ and let $x^{(-)}$ denote the string obtained from $x$ by deleting its last symbol. Note that $x=x^{(-)}a$ and $|x^{(-)}|\geq c$.
@@
We split $xy$ to two parts: a prefix $\hat{x}$ and a suffix $\hat{y}$ of $xy$ satisfying $\hat{x}\hat{y}=xy$. Let $\hat{a}$ denote the first symbol of $\hat{y}$ if $\hat{y}\neq\varepsilon$. Along the path $p_{i}$ from the root $r_{1}$ to $c_{s_{i}+1}$, $\rho_{1}(c_{s_{i}+1})=\hat{a}$ follows.